Circular No. 9045

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 K2 (CATALINA) An apparently asteroidal object discovered (as tabulated below) by

the Catalina Sky Survey (observer R. A. Kowalski), and posted on the Minor Planet Center’s ‘NEOCP’ confirmation page, has been found by CCD astrometrists to show cometary appearance. W. H. Ryan (Magdalena Ridge 2.4-m reflector, near Socorro, NM; May 18.4 UT) notes a distinct tail in p.a. ≈ 200^{o}. J. Young (Table Mountain 0.61-m reflector) reports a bright, well-defined coma of diameter 5′′ with a poorly defined 16′′ extension in p.a. ∼ 235^{o} on images taken during May 18.43–18.46 (light cirrus clouds present); he adds that additional images from May 19.33–19.35 show a small bright coma with a definite broad tail spanning p.a. 210^{o}–250^{o}, its greatest length being 14′′ in p.a. 235^{o}. L. Bittesini, D. Di Maria, E. Pettarin, and F. Piani (Farra D’Isonzo, Italy, 0.40-m reflector; May 18.9) find that forty co-added 60-s exposures reveal a 10′′ coma that is elongated in p.a. 190^{o}. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector; May 18.9) reports an almost-stellar center of diameter 4′′ with a 6′′ tail in p.a. 210^{o}. E. Guido, Castellammare di Stabia, Italy, writes that G. Sostero and he co-added 33 unfiltered 120-s exposures obtained remotely with a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., on May 19.2, revealing a tiny, compact coma ∼ 7′′ in diameter with a tail ∼ 12′′ long toward p.a. 240^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
May 18.33287 16^{h}39^{m}44.^{s}35 +60^{o}58′07′′.0 19.4 Kowalski

The available astrometry, the following very preliminary orbital elements, and an ephemeris appear on MPEC 2009-K20.

T = 2010 Aug. 18.870 TT  ω = 162.^{o}060
Ω = 112.706 }2000.0
q = 3.83854 AU  i = 70.910

COMET 220P/2009 H2 (McNAUGHT) J. L. Montani, Lunar and Planetary Laboratory, University of Arizona,

reports that he has examined the “incidental astrometry” images of this comet (cf. IAUC 9040) obtained on Apr. 28 with the Spacewatch 0.9-m Mosaic telescope at Kitt Peak by T. H. Bressi. In 1′′.1 seeing, the comet’s images appear to be stellar, with no trace of extended coma or tail, to the limiting magnitude of the ≈ 120-s single exposures.

2009 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green